# Complete asymptotics in the formation of quiescent big bang singularities

### v1

@@ 2.6. Estimating the frame
(73) $2t^{-2\varepsilon}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}\partial_{t}e_{I}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}=2t^{-2\varepsilon}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}(-Nk_{IJ}e_{J}^{i}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The expression $E_{\mathbf{I}}(Nk_{IJ}t^{-\varepsilon}e_{J}^{i})$ is a bad term. In particular, due to the arguments presented in Subsection 2.4, it can be written
@@
$2t^{-2\varepsilon}\textstyle{\int}_{\Sigma}k_{IJ}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}e_{J}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\geq-2(1+Ct^{\varepsilon})t^{-1}t^{-2\varepsilon}\textstyle{\sum}_{I,i}\textstyle{\int}_{\Sigma}|E_{\mathbf{I}}e_{I}^{i}|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
@@ 2.7. Estimating the second fundamental form and the structure coefficients
Moreover, if $\mathbf{I}=0$, then the only term we need to handle is the first expression in (80). The contribution to (79) from the first expression in (80) can be rewritten by integrating by parts; cf. (141). When doing so, there is one term involving the divergence of $e_{I}$. Due to (39), (54) and (60), this term can be estimated in absolute value by $C_{\ell}t^{-1+3\varepsilon}E_{\mathrm{tot},\ell}$. What remains to be considered is
@@ 2.8. The scalar field
At this point, we can appeal to (16). As a result, we need to estimate
@@ 2.10. Summing up
Combining Lemmas 29, 30 and 31 with (57) yields
@@ 3.1. Basic consequences of the assumptions
The conclusion is an immediate consequence of the assumptions and the argument given in the proof of [32, Lemma 75, p. 29]. ∎
@@
Note that $|(\ln t)(\mathring{p}_{I}-p_{I})|\leq C|\ln t|t^{\delta}$ is bounded for $t\in(0,T)$. Moreover, due to (36d), it is clear that $N^{-1}\leq 2$ and $(t\Theta)^{-1}\leq 2$. Due to this observation and the assumptions, it is clear that $N$, $N^{-1}$, $t\Theta$ and $(t\Theta)^{-1}$ are bounded. Combining the above observations, it is clear that
@@
It follows that $|E_{\mathrm{\mathbf{I}}}\ln\theta|\leq C$. Hence,
$\|\ln\theta\|_{C^{k_{0}}}\leq C\langle\ln t\rangle.$
Finally,
@@ 3.2. The eigenframe
Due to the assumptions, the eigenvalues $p_{I}$ of $\mathcal{K}$ are distinct everywhere. Due to [41, Lemma A.1, p. 223], we can, by taking a finite covering space of $\Sigma$, if necessary, assume that there are smooth vector fields $\widehat{e}_{I}$ on $M$, for $I=1,\ldots,n$, tangential to the hypersurfaces $\Sigma_{t}$ and such that $\mathcal{K}(\widehat{e}_{I})=p_{I}\widehat{e}_{I}$ and $h(\widehat{e}_{I},\widehat{e}_{J})=\delta_{IJ}$. In particular, $\{\widehat{e}_{I}\}_{I=1}^{n}$ is then an orthonormal frame for $h$ on each $\Sigma_{t}$. Denote the corresponding dual frame $\{\widehat{\omega}^{I}\}_{I=1}^{n}$. Next, we deduce evolution equations for the $\widehat{e}_{I}$.
@@
$\big|E_{\mathrm{\mathbf{I}}}A_{I}^{J}(s)-E_{\mathrm{\mathbf{I}}}A_{I}^{J}(r)\big|\leq\textstyle\int_{r}^{s}\big|E_{\mathrm{\mathbf{I}}}\partial_{t}A_{I}^{J}(t)\big|{\mathrm{d}}t\leq C(s^{2\delta}-r^{2\delta}).$
@@ 3.3. The expansion normalized eigenframe
The estimates for ${\check{e}}_{I}^{i}$ and ${\check{\omega}}_{i}^{I}$ follow directly from the definitions and Lemma 37, since $\langle\ln t\rangle^{k_{0}}t^{\delta}$ is bounded for $t\in(0,t_{0}]$. For $\Lambda_{IJ}^{K}$,
@@
The last estimate now follows from Lemma 37. ∎
@@ 3.4. Asymptotics
The result follows from the assumptions, the definition of $\mathcal{B}_{I}^{J}$, (30a), Lemmas 37 and 44, and Proposition 45. ∎
@@
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Let $0<\mathbf{c}_{1}\in\mathbb{R}$, $1\leq k\leq k_{0}-1$, $x\in\Sigma$, and suppose that the following estimates hold for some $m\in\mathbb{N}$, $t\leq t_{0}$ and all $|\mathrm{\mathbf{I}}|\leq k$: For each ${\check{e}}_{I}$ separately, either
@@
(133) $\begin{split}(\overline{\nabla}^{2}N^{\sharp})\big({\check{e}}_{I},{\check{\omega}}^{J}\big)&=\theta^{2p_{J}}(\overline{\nabla}^{2}N)({\check{e}}_{I},{\check{e}}_{J})\\ &=\theta^{2p_{J}}{\check{e}}_{I}{\check{e}}_{J}N+\theta^{2p_{J}}{\check{e}}_{I}(p_{J}\ln\theta){\check{e}}_{J}(N)+\theta^{2p_{J}}{\check{e}}_{J}(p_{I}\ln\theta){\check{e}}_{I}(N)\\ &\quad-\tfrac{1}{2}\theta^{2p_{J}}\Lambda_{IJ}^{K}{\check{e}}_{K}(N)+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{J}{\check{e}}_{K}(N)\\ &\quad+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}\Lambda_{JK}^{I}{\check{e}}_{K}(N).\end{split}$
@@
Let $3\leq n\in\mathbb{N}$, $k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Fix an open cover $U_{j}$ of $\Sigma$, with corresponding constants $\alpha_{j}$, as in Lemma 40, and let $m\in\mathbb{N}$. Suppose that for each $j$, the following pointwise estimates hold in $U_{j}$ for all $|\mathrm{\mathbf{I}}|\leq k$, $t\leq t_{0}$ and some $k\leq k_{0}-2$: for ${\check{e}}_{I}$,
@@
for all $t\leq t_{0}$. Finally, $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$ are $C^{\ell}$ initial data on the singularity.
@@
in $(0,t_{0}]\times U_{j}$, and for $|\mathrm{\mathbf{I}}|\leq\ell+1$, assuming $k_{0}\geq 2N+\ell+1$. By taking the maximum of the constants $C_{j}$, we get a constant $C$ such that
@@
for $|\mathrm{\mathbf{I}}|\leq\ell+1$, in $(0,t_{0}]\times\Sigma$.
@@
for $I\neq J$ and $|\mathrm{\mathbf{I}}|\leq\ell$. Combining this estimate with Lemmas 37, 49 and 51 yields
@@
Thus, by the assumptions and Lemmas 39 and 44,
@@ 4. Proof of the main result
At this point, we are ready to prove Theorem 11 by combining the main theorem of [32] with Theorems 18 and 22.
@@
Given admissibility thresholds $\sigma_{V},\sigma_{p}\in(0,1)$ as in the statement of the theorem, $\sigma$ is defined as in [32, (11), p. 6]:
@@
In [32, (12), p. 6], lower bounds on $k_{0}$ and $k_{1}$ are specified. Here we let
@@
i.e., $k_{0}$ is the minimal number such that [32, (12a), p. 6] holds. Note, however, that $k_{0}\geq 2$. We also assume that $k_{1}$ is the minimal integer such that [32, (12b), p. 6] holds. In particular, $\sigma$ is determined by $\sigma_{V}$ and $\sigma_{p}$; $k_{0}$ is determined by $n$; and $k_{1}$ is determined by $\sigma_{V}$, $\sigma_{p}$ and $n$. We define $\kappa:=k_{1}+2$.
@@
Next, note that the assumptions include the requirement that the eigenvalues of $\mathcal{K}_{0}$ be distinct. Due to the arguments presented in the proof of [41, Lemma A.1, p. 223], it follows that a finite (at most a $2^{n}$-fold) covering space of $\Sigma$ has a global frame. Going to this finite cover, we can assume that there is a global orthonormal frame $\{E_{i}\}_{i=1}^{n}$ of $(\Sigma,h_{\mathrm{ref}})$. Considering the Sobolev norms on the left hand side of (8), they are at most increased by a factor of $2^{n}$ when going to the finite cover. Moreover, the norms used in the statement of [32, Theorem 12, pp. 6-7] are defined using a frame $\{E_{i}\}_{i=1}^{n}$ and the norms in (8) are defined using the metric $h_{\mathrm{ref}}$. However, the two norms are equivalent up to constants depending only on $(\Sigma,h_{\mathrm{ref}})$ and the frame. In other words, this difference does not affect the statement. Fixing a $\zeta_{0}$ as in the statement of the theorem, there is a $\zeta_{0}$ as in the statement of [32, Theorem 12, pp. 6-7] (obtained by compensating for the covering space and the equivalence of the norms), say $\zeta_{0}^{\prime}$. Due to [32, Theorem 12, pp. 6-7], we then obtain a $\zeta_{1}$. Given that the assumptions of the theorem are satisfied with this $\zeta_{1}$ (and the original $\zeta_{0}$), the conclusions of [32, Theorem 12, pp. 6-7] hold. Moreover, the estimates derived in the proof of this theorem hold.
@@
Due to [32, Proposition 72, p. 29], there is a constant $\rho_{0}>0$ depending only on $\zeta_{0}$, $\sigma$, $k_{0}$, $k_{1}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In our setting, $\rho_{0}$ thus only depends on $\zeta_{0}$, $\sigma_{V}$, $\sigma_{p}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In accordance with [32, Notation 73, p. 29], we then say that $C$ is a standard constant if it is strictly positive and only depends on $\rho_{0}$, $\sigma_{p}$, $\sigma_{V}$, $k_{0}$, $k_{1}$, $c_{k_{1}+2}$ (see Definition 9), $(\Sigma,h_{\mathrm{ref}})$ and the orthonormal frame $\{E_{i}\}_{i=1}^{n}$.
@@
Next, we wish to appeal to Theorem 18. Note, therefore, that we have a solution to (11)–(19) as described in Subsection 1.4; the equations solved in [32] are the ones written down in [32, Proposition 68, pp. 23-24], and these equations coincide with (11)–(19). Next, note that (25) with $\varepsilon=\sigma$ is an immediate consequence of [32, (110), p. 37] and [32, Lemma 98, p. 39]; recall [32, Definition 85, p. 36]. The estimate (26) is an immediate consequence of [32, (147), p. 49]. Note also that $3\sigma\leq\sigma_{V}$ due to (135). This means that the conclusions of Theorem 18 hold.
@@
Next, we wish to combine the above conclusions with Theorem 22. From [32, Proposition 68, pp. 23-24], it follows that $g$ takes the form (29) and that (32) holds. Next, let $\mathring{\Phi}$ and $\mathring{\Psi}$ be as in the statement of Theorem 18 and let $\mathring{p}_{I}$ be the eigenvalues of $\mathring{\mathfrak{K}}$ (the matrix with components $\mathring{\mathfrak{K}}_{IJ}$). Then the $\mathring{p}_{I}$ are distinct due to [32, Theorem 12, pp. 6-7] and smooth since the $\mathring{\mathfrak{K}}_{IJ}$ are smooth. Due to [32, (14), p. 7], it is clear that (33) holds. Next, due to [32, (226), p. 70], we can, by demanding that $\zeta_{1}$ be large enough, ensure that $p_{I}+p_{J}-p_{K}\leq 1-\sigma_{p}/2$ for all $t\leq t_{0}$ and $I$, $J$, $K$ such that $I\neq J$. This means that (34) holds for any $\delta\leq\sigma$. Next, due to (28), the estimate (35) holds for all $k_{0}$, assuming $\delta<\sigma/4$. Assuming $\delta\leq\sigma/3$, the first three estimates in (36) follow immediately from (27). The final estimate follows from (27) by demanding that $T$ be small enough. By demanding that $T$ be small enough, we can also assume that $|p_{I}-p_{J}|>r$ for $I\neq J$ and some $r>0$. Fixing such $r$, $T$ and $\delta<\sigma/4$, it follows that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions for any $k_{0}\in\mathbb{N}$.
@@
At this point we can appeal to Theorem 22. This yields the conclusion that there are smooth data on the singularity, as desired. Moreover, (9) holds. The remaining conclusions of the theorem follow from [32, Theorem 12, pp. 6-7]. This finishes the proof of the theorem, with one caveat: we have demonstrated the desired conclusions for the finite covering space. On the other hand, the covering maps are local isometries, so that the conclusions descend to the quotient. The theorem follows. ∎
Appendix A Notation and conventions

### v2

@@ 2.6. Estimating the frame
(73) $2t^{-2\varepsilon}\textstyle{\sum}_{I,i}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}\partial_{t}e_{I}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}=2t^{-2\varepsilon}\textstyle{\sum}_{I,J,i}\textstyle{\int}_{\Sigma}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}(-Nk_{IJ}e_{J}^{i}){\mathrm{d}}\mu_{h_{\mathrm{ref}}}.$
The expression $Nk_{IJ}t^{-\varepsilon}e_{J}^{i}$ is a bad term. In particular, due to the arguments presented in Subsection 2.4, it can be written
@@
$2t^{-2\varepsilon}\textstyle{\sum}_{I,J,i}\textstyle{\int}_{\Sigma}k_{IJ}E_{\mathbf{I}}e_{I}^{i}\cdot E_{\mathbf{I}}e_{J}^{i}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}\geq-2(1+Ct^{\varepsilon})t^{-1}t^{-2\varepsilon}\textstyle{\sum}_{I,i}\textstyle{\int}_{\Sigma}|E_{\mathbf{I}}e_{I}^{i}|^{2}{\mathrm{d}}\mu_{h_{\mathrm{ref}}}$
@@ 2.7. Estimating the second fundamental form and the structure coefficients
Moreover, if $\mathbf{I}=0$, then the only term we need to handle is the first expression in (80). The contribution to (79) from the first expression in (80) can be rewritten by integrating by parts; cf. (141). When doing so, there is one term involving the divergence of $e_{I}$. Due to (39), this term can be estimated in absolute value by $C_{\ell}t^{-1+3\varepsilon}E_{\mathrm{tot},\ell}$. What remains to be considered is
@@ 2.8. The scalar field
Next, we appeal to (16). As a result, we need to estimate
@@ 2.10. Summing up
Combining Lemmas 31, 32 and 33 with (57) yields
@@ 3.1. Basic consequences of the assumptions
The conclusion is an immediate consequence of the assumptions and the argument given in the proof of [33, Lemma 77, p. 29]. ∎
@@
Note that $|(\ln t)(\mathring{p}_{I}-p_{I})|\leq C|\ln t|t^{\delta}$ is bounded for $t\in(0,T)$. Moreover, due to (36d), it is clear that $(t\Theta)^{-1}\leq 2$. Due to this observation and the assumptions, it is clear that $N$, $N^{-1}$, $t\Theta$ and $(t\Theta)^{-1}$ are bounded. Combining the above observations, it is clear that
@@
Thus $|E_{\mathrm{\mathbf{I}}}\ln\theta|\leq C$ and the second estimate in (116) follows. Finally,
@@ 3.2. The eigenframe
Due to the assumptions, the eigenvalues $p_{I}$ of $\mathcal{K}$ are distinct everywhere. Due to [42, Lemma A.1, p. 223], we can, by taking a finite covering space of $\Sigma$, if necessary, assume that there are smooth vector fields $\widehat{e}_{I}$ on $M$, for $I=1,\ldots,n$, tangential to the hypersurfaces $\Sigma_{t}$ and such that $\mathcal{K}(\widehat{e}_{I})=p_{I}\widehat{e}_{I}$ and $h(\widehat{e}_{I},\widehat{e}_{J})=\delta_{IJ}$. In particular, $\{\widehat{e}_{I}\}_{I=1}^{n}$ is then an orthonormal frame for $h$ on each $\Sigma_{t}$. Denote the corresponding dual frame $\{\widehat{\omega}^{I}\}_{I=1}^{n}$. Next, we deduce evolution equations for the $\widehat{e}_{I}$.
@@
$\big|E_{\mathrm{\mathbf{I}}}A_{I}^{J}(s)-E_{\mathrm{\mathbf{I}}}A_{I}^{J}(u)\big|\leq\textstyle\int_{u}^{s}\big|E_{\mathrm{\mathbf{I}}}\partial_{t}A_{I}^{J}(t)\big|{\mathrm{d}}t\leq C(s^{2\delta}-u^{2\delta}).$
@@ 3.3. The expansion normalized eigenframe
The estimates for ${\check{e}}_{I}^{i}$ and ${\check{\omega}}_{i}^{I}$ follow directly from the definitions, Lemma 39 and Lemma 49, since $\langle\ln t\rangle^{k_{0}-1}t^{\delta}$ is bounded for $t\in(0,t_{0}]$. For $\Lambda_{IJ}^{K}$,
@@
The last estimate now follows from Lemmas 39 and 49. ∎
@@ 3.4. Asymptotics
The result follows from the assumptions, the definition of $\mathcal{B}_{I}^{J}$, (30a), Lemmas 39 and 46, and Proposition 47. ∎
@@
Let $3\leq n\in\mathbb{N}$, $2\leq k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Let $0<\mathbf{c}_{1}\in\mathbb{R}$, $1\leq k\leq k_{0}-1$, $x\in\Sigma$, and suppose that the following estimates hold for some $m\in\mathbb{N}$, $t\leq t_{0}$ and all $|\mathrm{\mathbf{I}}|\leq k$: For each ${\check{e}}_{I}$ separately, either
@@
(133) $\begin{split}(\overline{\nabla}^{2}N)^{\sharp}\big({\check{e}}_{I},{\check{\omega}}^{J}\big)&=\theta^{2p_{J}}(\overline{\nabla}^{2}N)({\check{e}}_{I},{\check{e}}_{J})\\ &=\theta^{2p_{J}}{\check{e}}_{I}{\check{e}}_{J}N+\theta^{2p_{J}}{\check{e}}_{I}(p_{J}\ln\theta){\check{e}}_{J}(N)+\theta^{2p_{J}}{\check{e}}_{J}(p_{I}\ln\theta){\check{e}}_{I}(N)\\ &\quad-\tfrac{1}{2}\theta^{2p_{J}}\Lambda_{IJ}^{K}{\check{e}}_{K}(N)+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2p_{K}}\Lambda_{IK}^{J}{\check{e}}_{K}(N)\\ &\quad+\tfrac{1}{2}\textstyle\sum_{K}\theta^{2(p_{J}+p_{K}-p_{I})}\Lambda_{JK}^{I}{\check{e}}_{K}(N).\end{split}$
@@
Let $3\leq n\in\mathbb{N}$, $4\leq k_{0}\in\mathbb{N}$, $r>0$, $\delta\in(0,1)$ and fix a $\delta$-admissible potential $V\in C^{\infty}(\mathbb{R})$. Given that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions, assume that there is an eigenframe $\{\widehat{e}_{I}\}_{I=1}^{n}$ with the properties listed at the beginning of Subsection 3.2. Fix an open cover $U_{j}$ of $\Sigma$, with corresponding constants $\alpha_{j}$, as in Lemma 42, and let $m\in\mathbb{N}$. Suppose that for each $j$, the following pointwise estimates hold in $U_{j}$ for all $|\mathrm{\mathbf{I}}|\leq k$, $t\leq t_{0}$ and some $k\leq k_{0}-2$: for ${\check{e}}_{I}$,
@@
for all $t\leq t_{0}$. Finally, $(\Sigma,\mathring{\mathcal{H}},\mathring{\mathcal{K}},\mathring{\Phi},\mathring{\Psi})$ are $C^{\ell}$ robust nondegenerate quiescent initial data on the singularity for the Einstein–nonlinear scalar field equations.
@@
in $(0,t_{0}]\times U_{j}$, and for $|\mathrm{\mathbf{I}}|\leq\ell+2$, assuming $k_{0}\geq 2N+\ell+2$. By taking the maximum of the constants $C_{j}$, we get a constant $C$ such that
@@
for $|\mathrm{\mathbf{I}}|\leq\ell+2$, in $(0,t_{0}]\times\Sigma$.
@@
for $I\neq J$ and $|\mathrm{\mathbf{I}}|\leq\ell+1$. Combining this estimate with Lemmas 39, 51 and 53 yields
@@
Thus, by the assumptions and Lemmas 41 and 46,
@@ 4. Proof of the main result
At this point, we are ready to prove Theorem 11 by combining the main theorem of [33] with Theorems 19 and 24.
@@
Given admissibility thresholds $\sigma_{V},\sigma_{p}\in(0,1)$ as in the statement of the theorem, $\sigma$ is defined as in [33, (11), p. 6]:
@@
In [33, (12), p. 6], lower bounds on $k_{0}$ and $k_{1}$ are specified. Here we let
@@
i.e., $k_{0}$ is the minimal number such that [33, (12a), p. 6] holds. Note, however, that $k_{0}\geq 2$. We also assume that $k_{1}$ is the minimal integer such that [33, (12b), p. 6] holds. In particular, $\sigma$ is determined by $\sigma_{V}$ and $\sigma_{p}$; $k_{0}$ is determined by $n$; and $k_{1}$ is determined by $\sigma_{V}$, $\sigma_{p}$ and $n$. We define $\kappa:=k_{1}+2$.
@@
Next, note that the assumptions include the requirement that the eigenvalues of $\mathcal{K}_{0}$ be distinct. Due to the arguments presented in the proof of [42, Lemma A.1, p. 223], it follows that a finite (at most a $2^{n}$-fold) covering space of $\Sigma$ has a global frame. Going to this finite cover, we can assume that there is a global orthonormal frame $\{E_{i}\}_{i=1}^{n}$ of $(\Sigma,h_{\mathrm{ref}})$. Considering the Sobolev norms on the left hand side of (8), they are at most increased by a factor of $2^{n}$ when going to the finite cover. Moreover, the norms used in the statement of [33, Theorem 12, pp. 6-7] are defined using a frame $\{E_{i}\}_{i=1}^{n}$ and the norms in (8) are defined using the metric $h_{\mathrm{ref}}$. However, the two norms are equivalent up to constants depending only on $(\Sigma,h_{\mathrm{ref}})$ and the frame. In other words, this difference does not affect the statement. Fixing a $\zeta_{0}$ as in the statement of the theorem, there is a $\zeta_{0}$ as in the statement of [33, Theorem 12, pp. 6-7] (obtained by compensating for the covering space and the equivalence of the norms), say $\zeta_{0}^{\prime}$. Due to [33, Theorem 12, pp. 6-7], we then obtain a $\zeta_{1}$. Given that the assumptions of the theorem are satisfied with this $\zeta_{1}$ (and the original $\zeta_{0}$), the conclusions of [33, Theorem 12, pp. 6-7] hold. Moreover, the estimates derived in the proof of this theorem hold.
@@
Due to [33, Proposition 74, p. 29], there is a constant $\rho_{0}>0$ depending only on $\zeta_{0}$, $\sigma$, $k_{0}$, $k_{1}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In our setting, $\rho_{0}$ thus only depends on $\zeta_{0}$, $\sigma_{V}$, $\sigma_{p}$, $(\Sigma,h_{\mathrm{ref}})$ and $\{E_{i}\}_{i=1}^{n}$. In accordance with [33, Notation 75, p. 29], we then say that $C$ is a standard constant if it is strictly positive and only depends on $\rho_{0}$, $\sigma_{p}$, $\sigma_{V}$, $k_{0}$, $k_{1}$, $c_{k_{1}+2}$ (see Definition 9), $(\Sigma,h_{\mathrm{ref}})$ and the orthonormal frame $\{E_{i}\}_{i=1}^{n}$.
@@
Next, we wish to appeal to Theorem 19. Note, therefore, that we have a solution to (11)–(19) as described in Subsection 1.4; the equations solved in [33] are the ones written down in [33, Proposition 70, pp. 23-24], and these equations coincide with (11)–(19). Next, note that (25) with $\varepsilon=\sigma$ is an immediate consequence of [33, (110), p. 37] and [33, Lemma 100, p. 39]; recall [33, Definition 87, p. 36]. The estimate (26) is an immediate consequence of [33, (147), p. 49]. Note also that $3\sigma\leq\sigma_{V}$ due to (135). This means that the conclusions of Theorem 19 hold.
@@
Next, we wish to combine the above conclusions with Theorem 24. From [33, Proposition 70, pp. 23-24], it follows that $g$ takes the form (29) and that (32) holds. Next, let $\mathring{\Phi}$ and $\mathring{\Psi}$ be as in the statement of Theorem 19 and let $\mathring{p}_{I}$ be the eigenvalues of $\mathring{\mathfrak{K}}$ (the matrix with components $\mathring{\mathfrak{K}}_{IJ}$). Then the $\mathring{p}_{I}$ are distinct due to [33, Theorem 12, pp. 6-7] and smooth since the $\mathring{\mathfrak{K}}_{IJ}$ are smooth. Due to [33, (14), p. 7], it is clear that (33) holds. Next, due to [33, (226), p. 70], we can, by demanding that $\zeta_{1}$ be large enough, ensure that $p_{I}+p_{J}-p_{K}\leq 1-\sigma_{p}/2$ for all $t\leq t_{0}$ and $I$, $J$, $K$ such that $I\neq J$. This means that (34) holds for any $\delta\leq\sigma$. Next, due to (28), the estimate (35) holds for all $k_{0}$, assuming $\delta<\sigma/4$. Assuming $\delta\leq\sigma/3$, the first three estimates in (36) follow immediately from (27). The final estimate follows from (27) by demanding that $T$ be small enough. By demanding that $T$ be small enough, we can also assume that $|p_{I}-p_{J}|>r$ for $I\neq J$ and some $r>0$. Fixing such $r$, $T$ and $\delta<\sigma/4$, it follows that $(M,g,\varphi)$ is a solution to the Einstein–nonlinear scalar field equations satisfying the $(\delta,n,k_{0},r)$-assumptions for any $k_{0}\in\mathbb{N}$.
@@
At this point we can appeal to Theorem 24. This yields the conclusion that there are smooth data on the singularity, as desired. Moreover, (9) holds. The remaining conclusions of the theorem follow from [33, Theorem 12, pp. 6-7]. This finishes the proof of the theorem, with one caveat: we have demonstrated the desired conclusions for the finite covering space. On the other hand, the covering maps are local isometries, so that the conclusions descend to the quotient. The theorem follows. ∎
Appendix A Notation, conventions, estimates and identities